OpenStudy (anonymous):

given that the x-intercept of a line is (-2 1/2,0) and y-intercept is (0,5), what is the equation of the line in slope-intercept form

OpenStudy (anonymous):

Its simple friend... slope m=(y2-y2)/(x2-x1) then y-y1=m(x-x1) put x1=-2.5 y1=0 x2=0 y2=5 then solve it....

OpenStudy (anonymous):

Is it clear are not...?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Answer is y=2x+5